# Bayesian Inference Under Differential Privacy: Prior Selection Considerations with Application to Univariate Gaussian Data and Regression

### v1

@@ 3 Incorporating Constraints
By incorporating constraints on data values, analysts can reduce the uncertainty in estimates of the parameters of interest “for free.” They need not make additional assumptions; they merely determine how the conditions on the data values constrain the model used for analysis.
3.1 Constraints in the Univariate Gaussian Setting
@@
In the univariate Gaussian setting from Section 2.2, the constraint that $Y_{i}\in[0,1]$ affects the model in two ways. Firstly, the bound constrains the parameters $\mu$ and $\sigma^{2}$, since bounded random variables have bounded expectations and variances. Secondly, the bound constrains the sufficient statistics $\bar{Y}$ and $S^{2}$, since samples of bounded random variables have bounded means and variances. We demonstrate how we establish these bounds in a manner that can be easily incorporated into the Gibbs sampler from Section 2.2. See Appendix E for proofs.
@@
First, Theorem 2 establishes conditional bounds for $\mu\mid\sigma^{2}$ and $\sigma^{2}\mid\mu$ that can be incorporated into the prior. That is, since the data values are bounded, we may replace the unbounded normal-inverse gamma prior from (2) with a prior of the same form truncated to be within the feasible region. This yields a posterior of the same form as in Section 2.2 truncated to the feasible region.
@@
Let $Y_{1},\ldots,Y_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$ and suppose each $Y_{i}\in[0,1]$. If $\mu$ is known, then $\sigma^{2}\in[0,\mu(1-\mu)]$. If $\sigma^{2}$ is known, then $\mu\in[1/2-\sqrt{1/4-\sigma^{2}},1/2+\sqrt{1/4-\sigma^{2}}]$.
@@
If $Y_{1},\ldots,Y_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$ and $Y_{i}\in[0,1]$, then $\sigma^{2}\leq 1/4$.
Theorem 3 establishes bounds for $\bar{Y}\mid S^{2}$ and $S^{2}\mid\bar{Y}$ that can be incorporated into the likelihood. This yields a posterior of the same form as in Section 2.2 truncated to the feasible region.
@@
Let $Y_{1},\ldots,Y_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$ and suppose each $Y_{i}\in[0,1]$. If $\bar{Y}$ is known, then $S^{2}\in[0,n/(n-1)\bar{Y}(1-\bar{Y})]$. If $S^{2}$ is known, then $\bar{Y}\in[1/2-\sqrt{1/4-(n-1)/n\cdot S^{2}},1/2+\sqrt{1/4-(n-1)/n\cdot S^{2}}]$.
@@
3.2 Example: Constraints for the Blood Lead Levels Dataset
We now demonstrate the effect of incorporating these constraints using genuine data. [^2] Code to reproduce all experiments can be found at https://github.com/zekicankazan/dp_priors All computations are performed on the $[0,1]$ scale, but all quantities are converted back to the original scale for clarity of presentation. Interval estimates are those with highest posterior density (HPD). [^3] HPD intervals are preferred over quantile-based intervals when the posterior distribution is skewed [27]. HPD intervals can be computed with standard tools, such as the HDInterval package in R [30].
@@ 3.2 Example: Constraints for the Blood Lead Levels Dataset
Researchers sampled the blood lead levels of $n=43$ policemen assigned to outdoor work in Egypt to examine the effect of leaded gasoline on exposed individuals [32]. The sample mean was $\bar{Y}=32.08\,\mu g/dL$, and the sample variance was $S^{2}=16.98^{2}\,\mu g^{2}/dL^{2}$. An expert advises that blood lead levels in this region are reasonably bounded above by $100\,\mu g/dL$. A realization of the Laplace Mechanism with $\varepsilon_{1}=\varepsilon_{2}=0.25$ yields noisy statistics $\bar{Y}^{*}=34.30\,\mu g/dL$ and $S^{2*}=47.16^{2}\,\mu g^{2}/dL^{2}$. An analyst wishes to estimate $\mu$, the average blood lead level of all policemen assigned to outdoor work in Egypt. She believes blood lead levels are reasonably approximated by a Gaussian distribution. To reflect prior knowledge, she uses (2) with $(\mu_{0}=12.5,\sigma_{0}^{2}=3.8^{2},\kappa_{0}=1,\nu_{0}=1)$, roughly representing the equivalent information of a single “prior data point.”
Figure 1: Joint posterior distribution for $(\mu,\sigma^{2})$ in Example 1. Here, $(\bar{Y}^{*}=34.30,S^{2*}=47.17^{2})$ is represented by the red circle, the unreleased $(\bar{Y}=32.08,S^{2}=16.96^{2})$ is represented by the green triangle, the analyst’s posterior mode is represented by the blue diamond, and $(\mu_{0}=12.5,S^{2}=3.8^{2})$ is represented by the purple square. Upper and lower panels display the posterior when constraints are and are not accounted for, respectively. The shaded area represents the feasible region for $(\mu,\sigma^{2})$ from Theorem 2. The main and secondary axes present the original and $[0,1]$ scale, respectively. Plots based on 5,000 Gibbs sampler iterations.
@@
We emphasize that the constrained posterior is not merely a truncated version of the unconstrained posterior; truncation in this way would yield inaccurate point and interval estimates. Instead, by truncating within the Gibbs sampler, we have mathematically coherent estimates.
@@ 4 Default Prior Choices
As evident in Example 1, the results of a Bayesian analysis under DP can be sensitive to the choice of prior distribution. This is unsurprising, since the addition of noise yields data with less information about model parameters than in the public setting, making it difficult to overwhelm the information in the prior distribution. One potential fix—used in examples in the Bayesian DP inference literature—is to increase the prior variance, yielding a weakly informative prior. One also could examine the limit in which the prior variance is infinite. We show, however, that this strategy may lead to improper limiting posterior distributions, even in settings where the nonprivate posterior is proper. In such cases, weakly informative priors may yield unreliable inference and should be avoided [14].
4.1 Default Priors in the Univariate Gaussian Setting
For the unconstrained analysis in the univariate Gaussian setting, a weakly informative prior of the form in (2) is created by making $\kappa_{0}$ and $\nu_{0}$ very small. In the limit as these hyperparameters go to zero, this produces the default prior $p(\mu,\sigma^{2})\propto(\sigma^{2})^{-1}$. In the nonprivate setting, this is called the independent Jeffrey’s prior [38]. While it does not correspond to a proper probability distribution, the posterior distribution it produces is not only proper, but also frequentist matching. [^4] In general, Bayesian credible intervals do not have the property that a 95% interval will cover the true value 95% of the time under repeated sampling. Frequentist matching priors, however, produce credible intervals identical to corresponding frequentist confidence intervals and so 95% intervals have 95% coverage (see [8]). In the private setting, unfortunately, the posterior distribution produced by $p(\mu,\sigma^{2})\propto(\sigma^{2})^{-1}$ is not a proper probability distribution. The following result, proved in Appendix F, demonstrates this.
@@ 4.1 Default Priors in the Univariate Gaussian Setting
For confidential data $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}{\mathcal{N}}(\mu,\sigma^{2})$ where sufficient statistics $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$ are released, if an analyst has prior $p(\mu,\sigma^{2})\propto(\sigma^{2})^{-1}$, then their posterior $p(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ is not a proper probability distribution.
@@
For confidential data $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}{\mathcal{N}}(\mu,\sigma^{2})$ where $n>3$ and sufficient statistics $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$ are released, if an analyst has prior $p(\mu,\sigma^{2})\propto 1$, then their posterior $p(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ is a proper probability distribution.
@@
Figure 2: The empirical coverage rate (top) and average length (bottom) of 95% HPD intervals for $\mu$ for different sample sizes. Results based on 10,000 simulated datasets $Y_{i}\in[0,1]$ released with $\varepsilon_{1}=\varepsilon_{2}=0.1$ and analyzed with prior $p(\mu,\sigma^{2})\propto 1$. Analyses with constraints accounted for are solid lines and not accounted for are dashed lines. Data generating model is either ${\mathcal{N}}(\mu=0.1,\sigma^{2}=0.04^{2})$ (red) or ${\mathcal{N}}(\mu=0.5,\sigma^{2}=0.2^{2})$ (blue). Each Gibbs sampler is run for 20,000 iterations.
@@
We use simulation experiments to assess the properties of the posterior distributions based on priors that use or disregard the constraints on the data values. For each of 10,000 simulated datasets of a given size $n$, we create a 95% HPD interval based on the Gibbs sampler in Appendix D. Figure 2 displays the results; see Appendix J for run-time details. Without enforcing constraints, the 95% credible interval has approximately the nominal 95% coverage rate regardless of $n$, and the average interval length is linearly related to $n$ on the log-log scale. [^5] A regression of log average interval length on log $n$ has a slope of $-1$ (with $R^{2}>0.999$), indicating that average interval length decays proportionally to $1/n$. A similar result holds for the RMSE; see Appendix H. When enforcing the constraints, the results are similar to those for the unconstrained analysis when $n$ is relatively large. In these cases, the credible intervals for $\mu$ are far from the boundary, so that the constraints have little effect. When $n$ is relatively small, the 95% credible interval has lower than 95% coverage rate when the true $\mu$ is close to the boundary and near 100% coverage when $\mu$ is in the middle of the range. In fact, it can be shown that the prior $p(\mu,\sigma^{2})\propto 1$ over the region where $\mu\in[0,1]$ and $\sigma^{2}\in(0,\mu(1-\mu)]$ induces the marginal distribution $\mu\sim\textsf{Beta}(2,2)$. This Beta distribution places more probability density in the center of the distribution than near 00 or $1$, resulting in the observed over and under coverage, depending on the true value of $\mu$. This suggests that $p(\mu,\sigma^{2})\propto 1$ is most appropriate when an analyst believes $\mu$ more likely is near the center of the distribution than the tails. That is, while the prior is “non-informative” in the unconstrained analysis, it is not so in the constrained analysis.
@@
Enforcing the constraint, however, does have advantages. For all points in Figure 2, the average interval length for the constrained analysis is less than $0.75$. Meanwhile, the average interval length without constraints is greater than $1$ for all $n\leq 50$, indicating that the credible intervals must include values that are not within $[0,1]$. Such intervals often include all of $[0,1]$, yielding intervals that are useless in practice. Indeed, the calibrated coverage rate for the unconstrained analysis is of dubious merit, as these credible intervals are inflated by permitting probability mass on infeasible regions of support. Additionally, when we estimate $\mu$ with its posterior mode, the RMSEs are uniformly lower in the constrained analysis than in the unconstrained; see Figure 8 in Appendix H. When $\mu=0.5$ the constrained analysis delivers a “win-win” result: it has a higher coverage rate with a shorter average interval length and a lower RMSE than the unconstrained analysis.
@@
4.2 Default Priors for the Blood Lead Levels Example
@@
Figure 3: Plot of the joint posterior distribution for $(\mu,\sigma^{2})$ in Example 1 under a uniform prior. The point $(\bar{Y}^{*}=34.30,S^{2}=47.17^{2})$ is represented by the red circle, the unreleased point $(\bar{Y}=32.08,S^{2}=16.96^{2})$ is represented by the green triangle, and the analyst’s posterior mode is represented by the blue diamond. The upper and lower panels provide the posterior when constraints are and are not accounted for, respectively. The shaded area represents the feasible region for $(\mu,\sigma^{2})$ from Theorem 2. The main axes and secondary axes present the original and $[0,1]$ scale, respectively. This plot is based on 5,000 Gibbs iterations.
@@
Analysts also might be interested in using the Bayesian model for prediction. Figure 4 plots the posterior predictive distribution for the uniform prior, which is computed via $Y_{\textup{new}}^{(t)}\sim{\mathcal{N}}(\mu^{(t)},\sigma^{2(t)})$ for each Gibbs iteration $t$, under the unconstrained and constrained analyses. [^6] For the constrained analysis, the distribution is replaced by a truncated Gaussian. The constrained posterior predictive distribution only includes values in the feasible region of 00 to $100\,\mu\textup{g}/\textup{dL}$, with the distribution centered around the posterior mode. The unconstrained posterior predictive distribution, meanwhile, generates 24% of predictive draws as negative values and 10% as values greater than $100\,\mu\textup{g}/\textup{dL}$. To obtain reasonable predictions with the draws from the unconstrained analysis, an analyst must resort to ad hoc methods. They might, for example, re-code all negative predictions to 00 and all predictions greater than $100$ to $100$. This would, among other effects, lead to 24% of predictions being exactly zero, which is not reasonable scientifically. Alternatively, the analyst could sample predictions from a truncated Gaussian distribution. The variance of this distribution is still substantially inflated by the samples of overly large values of $\sigma^{2(t)}$. Thus, predictions have larger variability under this ad hoc approach than under the theoretically principled constrained analysis.
@@
Figure 4: Posterior predictive distribution of a new observation given the posterior draws from Example 1 and prior $p(\mu,\sigma^{2})\propto 1$. The value $\bar{Y}^{*}=34.30$ is represented by the solid red line, the (unreleased) value $\bar{Y}=32.08$ is represented by the dotted green line, and the analyst’s posterior mode is represented by the blue dot-dashed line. Upper and lower panels provide the posterior predictive distributions without and without accounting for constraints, respectively. The shaded area represents the feasible region for a new observation. The main axis and secondary axis present the original and $[0,1]$ scale, respectively. Plots based on 100,000 Gibbs sampler iterations.
5 Non-Univariate Gaussian Application

### v2

@@ 3 ENFORCING CONSTRAINTS
The modeling in Section 2.2 does not account for the fact that the data values lie within some $[a,b]$. In this section, we show that analysts who incorporate the constraints on the data values can reduce the uncertainty in estimates of parameters of interest “for free.” They need not make additional assumptions; they merely determine how conditions on the data values constrain the model used for analysis.
There are several ways an analyst could incorporate these constraints. In this work, we introduce implied constraints on all unknown parameters as part of the modeling and sampling. This leads to straightforward estimation methods that can be readily generalized beyond the simple normal model. Another approach is to put constraints directly on the likelihood function for the model. We illustrate this approach for the normal model in Appendix I. Ultimately, we find that our approach is computationally advantageous and can offer more reliable repeated sampling performance; see Appendix I for details.
3.1 Constraints for the Gaussian Setting
@@
In the univariate Gaussian setting from Section 2.2, the constraint that $Y_{i}\in[0,1]$ affects the model in two ways. Firstly, the bound constrains the parameters $\mu$ and $\sigma^{2}$, since bounded random variables have bounded moments. Secondly, the bound constrains the sufficient statistics $\bar{Y}$ and $S^{2}$, since samples of bounded random variables have bounded means and variances. We demonstrate how we establish these bounds in a manner that can be easily incorporated into the Gibbs sampler from Section 2.2. See Appendix E for proofs.
@@
First, Theorem 2 establishes conditional bounds for $\mu\mid\sigma^{2}$ and $\sigma^{2}\mid\mu$. These bounds follow from monotonicity of expectation. As a result, since the data values are assumed to be bounded, we may replace the unbounded normal-inverse gamma prior from (2) with a prior of the same form truncated to be within the feasible region. This yields a posterior of the same form as in Section 2.2 truncated to the feasible region.
@@
Let $Y_{i}\in[0,1]$ have moments $E[Y_{i}]=\mu$ and $V[Y_{i}]=\sigma^{2}$. It follows that $\sigma^{2}\in[0,\mu(1-\mu)]$ and $\mu\in[1/2-\sqrt{1/4-\sigma^{2}},1/2+\sqrt{1/4-\sigma^{2}}]$.
@@
If $Y_{i}\in[0,1]$, then $V[Y_{i}]=\sigma^{2}\leq 1/4$.
Theorem 3 establishes conditional bounds for $\bar{Y}\mid S^{2}$ and $S^{2}\mid\bar{Y}$. This yields a posterior of the same form as in Section 2.2 truncated to the feasible region.
@@
Let $Y_{1},\ldots,Y_{n}$ be such that each $Y_{i}\in[0,1]$. Then $S^{2}\in[0,n/(n-1)\bar{Y}(1-\bar{Y})]$ and $\bar{Y}\in[1/2-\sqrt{1/4-(n-1)/n\cdot S^{2}},1/2+\sqrt{1/4-(n-1)/n\cdot S^{2}}]$.
@@
3.2 Example: The Blood Lead Dataset
We now demonstrate the effect of incorporating these constraints using genuine data. [^2] Code to reproduce all experiments is in Appendix L and at https://github.com/zekicankazan/dp_priors. All computations are performed on the $[0,1]$ scale, but all quantities are converted back to the original scale for clarity of presentation. Interval estimates are those with highest posterior density (HPD). [^3] HPD intervals are preferred over quantile-based intervals when the posterior distribution is skewed (Kruschke, 2015). HPD intervals can be computed with standard tools, such as the HDInterval package in R (Meredith & Kruschke, 2020, License: GPL-3). The following example, adapted from an introductory statistics textbook (Diez et al., 2012, Section 7.1), is representative of a common statistical analysis for numerical, scalar data.
@@ 3.2 Example: The Blood Lead Dataset
Researchers sampled the blood lead levels of $n=43$ policemen assigned to outdoor work in Egypt to examine the effect of leaded gasoline on exposed individuals (Mortada et al., 2001). The sample mean was $\bar{Y}=32.08\,\mu g/dL$, and the sample variance was $S^{2}=16.98^{2}\,\mu g^{2}/dL^{2}$. An expert advises that blood lead levels in this region are reasonably bounded above by $100\,\mu g/dL$. Using these bounds, the researchers use the Laplace Mechanism with $\varepsilon_{1}=\varepsilon_{2}=0.25$ to release noisy statistics $\bar{Y}^{*}=34.30\,\mu g/dL$ and $S^{2*}=47.16^{2}\,\mu g^{2}/dL^{2}$. A secondary data analyst seeks inferences for $\mu$ and $\sigma^{2}$, the average and variance of blood lead levels of all policemen assigned to outdoor work in Egypt. She believes blood lead levels are reasonably approximated by a Gaussian distribution. To reflect prior knowledge, she uses (2) with $(\mu_{0}=12.5,\sigma_{0}^{2}=3.8^{2},\kappa_{0}=1,\nu_{0}=1)$, roughly representing the equivalent information of a single “prior data point.”
Figure 1: Joint posterior distribution for $(\mu,\sigma^{2})$ in Example 1. Here, $(\bar{Y}^{*}=34.30,S^{2*}=47.17^{2})$ is represented by the red circle, the unreleased $(\bar{Y}=32.08,S^{2}=16.96^{2})$ is represented by the green triangle, the analyst’s posterior mode is represented by the blue diamond, and $(\mu_{0}=12.5,S^{2}=3.8^{2})$ is represented by the purple square. Upper and lower panels display the posterior when constraints are and are not accounted for, respectively. The shaded area represents the feasible region for $(\mu,\sigma^{2})$ from Theorem 2. Plots based on 5,000 Gibbs sampler iterations.
@@
The constrained posterior is not merely a truncated version of the unconstrained posterior; truncation in this way would yield inaccurate point and interval estimates. Instead, by truncating within the Gibbs sampler, we have mathematically coherent estimates.
@@ 4 DEFAULT PRIOR CHOICES
As evident in Example 1, the results of a Bayesian analysis under DP can be sensitive to the prior distribution. This is unsurprising, since the addition of noise yields data with less information about model parameters than in the public setting, making it difficult to overwhelm the information in the prior distribution. One potential fix—used in examples in the Bayesian DP inference literature—is to increase the prior variance, yielding a weakly informative prior. One also could examine the limit in which the prior variance is infinite. We show, however, that this strategy may lead to improper limiting posterior distributions, even in settings where the nonprivate posterior is proper. In such cases, weakly informative priors may yield unreliable inference and should be avoided (Gelman, 2006).
4.1 Default Priors for the Gaussian Setting
For the unconstrained analysis in the univariate Gaussian setting, a weakly informative prior of the form in (2) is created by making $\kappa_{0}$ and $\nu_{0}$ small. In the limit as these hyperparameters go to zero, this produces the default prior $p(\mu,\sigma^{2})\propto(\sigma^{2})^{-1}$. In the nonprivate setting, this is called the independent Jeffrey’s prior (Sun & Berger, 2007). While it does not correspond to a proper probability distribution, the posterior distribution it produces is not only proper, but also frequentist matching. In the private setting, unfortunately, the posterior distribution produced by $p(\mu,\sigma^{2})\propto(\sigma^{2})^{-1}$ is not a proper probability distribution. The following result, proved in Appendix F, demonstrates this.
@@ 4.1 Default Priors for the Gaussian Setting
For confidential data $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}{\mathcal{N}}(\mu,\sigma^{2})$ where $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$ are released, if an analyst has prior $p(\mu,\sigma^{2})\propto(\sigma^{2})^{-1}$, then their posterior $p(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ is not a proper probability distribution.
@@
Figure 2: The average length (top) and coverage rate (bottom) of 95% HPD intervals for $\mu$ for different $n$. Results based on 10,000 simulated datasets $Y_{i}\in[0,1]$ released with $\varepsilon_{1}=\varepsilon_{2}=0.1$ and analyzed with prior $p(\mu,\sigma^{2})\propto 1$. Analyses with constraints accounted for are solid lines and not accounted for are dashed lines. Data generating model is either ${\mathcal{N}}(\mu=0.1,\sigma^{2}=0.04^{2})$ (red) or ${\mathcal{N}}(\mu=0.5,\sigma^{2}=0.2^{2})$ (blue). Note that the red dashed line is below the blue dashed line. Each Gibbs sampler is run for 20,000 iterations.
@@
For confidential data $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}{\mathcal{N}}(\mu,\sigma^{2})$ where $n>3$ and $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$ are released, if an analyst has prior $p(\mu,\sigma^{2})\propto 1$, then their posterior $p(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ is a proper probability distribution.
@@
We use simulation experiments to assess the properties of the posterior distributions based on priors that use or disregard the constraints on the data values. For each of 10,000 simulated datasets of a given $n$, we create a 95% HPD interval based on the Gibbs sampler in Appendix D. Figure 2 displays the results; see Appendix K for run-time details. Without enforcing constraints, the 95% credible interval has approximately the nominal 95% coverage rate regardless of $n$, and the average interval length decays proportionally to $1/n$.  [^4] A regression of log average interval length on log $n$ has a slope of $-1$ (with $R^{2}>0.999$), indicating that average interval length decays proportionally to $1/n$. A similar result holds for the RMSE; see Appendix H. When enforcing the constraints, the results are similar to those for the unconstrained analysis when $n$ is large. In these cases, credible intervals for $\mu$ are far from the boundary, so that the constraints have little effect. When $n$ is small, the 95% credible interval has lower than 95% coverage rate when the true $\mu$ is close to the boundary and near 100% coverage when $\mu$ is in the middle of the range. In fact, it can be shown that the prior $p(\mu,\sigma^{2})\propto 1$ over the region where $\mu\in[0,1]$ and $\sigma^{2}\in(0,\mu(1-\mu)]$ induces the marginal distribution $\mu\sim\textsf{Beta}(2,2)$. This Beta distribution places more probability density in the center of the distribution than near 00 or $1$, resulting in the observed over and under coverage, depending on the true value of $\mu$. This suggests that $p(\mu,\sigma^{2})\propto 1$ is most appropriate when an analyst believes $\mu$ more likely is near the center of the distribution than the tails. That is, while the prior is “non-informative” in the unconstrained analysis, it is not so in the constrained analysis.
@@
Enforcing the constraint, however, does have advantages. For all points in Figure 2, the average interval length for the constrained analysis is less than $0.75$. Meanwhile, the average interval length without constraints is greater than $1$ for $n\leq 50$, indicating that the credible intervals must include values that are not within $[0,1]$. Such intervals often include all of $[0,1]$; these are practically useless. Indeed, the calibrated coverage rate for the unconstrained analysis is of dubious merit, as these credible intervals are inflated by permitting probability mass on infeasible regions. Additionally, when we estimate $\mu$ with its posterior mode, the RMSE is uniformly lower in the constrained analysis than in the unconstrained; see Figure 8 in Appendix H. In practical terms, the incorporation of constraints yields estimates that are more sensible scientifically, have less error, and have tighter interval estimates.
@@
4.2 Default Priors for Example 1
Figure 3: Plot of the joint posterior distribution for $(\mu,\sigma^{2})$ in Example 1 under a uniform prior. The point $(\bar{Y}^{*}=34.30,S^{2}=47.17^{2})$ is represented by the red circle, the unreleased point $(\bar{Y}=32.08,S^{2}=16.96^{2})$ is represented by the green triangle, and the analyst’s posterior mode is represented by the blue diamond. The upper and lower panels provide the posterior when constraints are and are not accounted for, respectively. The shaded area represents the feasible region for $(\mu,\sigma^{2})$ from Theorem 2. This plot is based on 5,000 Gibbs iterations.
@@ 4.2 Default Priors for Example 1
Analysts also might be interested in using the Bayesian model for prediction. Figure 4 plots the posterior predictive distribution for the uniform prior, which is computed via $Y_{\textup{new}(t)}\sim{\mathcal{N}}(\mu_{(t)},\sigma^{2}_{(t)})$ for each Gibbs iteration $t$, under the unconstrained and constrained analyses. [^5] For the constrained analysis, the distribution is replaced by a truncated Gaussian. The constrained posterior predictive distribution only includes values in the feasible region of 00 to $100\,\mu\textup{g}/\textup{dL}$, with the distribution centered around the posterior mode. The unconstrained posterior predictive distribution, meanwhile, generates 24% of predictive draws as negative values and 10% as values greater than $100\,\mu\textup{g}/\textup{dL}$. This leads to the constrained analysis having substantially lower predictive variance than the unconstrained: the standard deviation is $53\,\mu\textup{g}/\textup{dL}$ without constraints and $25\mu\textup{g}/\textup{dL}$ with constraints.
To obtain reasonable predictions with the draws from the unconstrained analysis, an analyst must resort to ad hoc methods. They might, for example, re-code all negative predictions to 00 and all predictions greater than $100$ to $100$. This would lead to 24% of predictions being exactly zero, which is not reasonable scientifically, and results in an inflated predictive standard deviation of $35\,\mu\textup{g}/\textup{dL}$. Alternatively, the analyst could sample predictions from a truncated Gaussian distribution. The variance of this distribution is still inflated by the samples of overly large values of $\sigma^{2}_{(t)}$, leading to a predictive standard deviation of $27\,\mu\textup{g}/\textup{dL}$ under this ad hoc approach, which is larger than that of the theoretically principled constrained analysis.
@@
Figure 4: Posterior predictive distribution of a new observation for the posterior draws from Example 1 and prior $p(\mu,\sigma^{2})\propto 1$. Upper and lower panels provide the posterior predictive distributions without and without accounting for constraints, respectively. The shaded area represents the feasible region for a new observation. Plots based on 100,000 Gibbs sampler iterations.
5 REGRESSION APPLICATION